Circular No. 9134

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

CBAT Director: Daniel W. E. Green; Room 209; Dept. of Earth and Planetary Sciences; Harvard University; 20 Oxford St.; Cambridge, MA 02138, U.S.A.

cbat@iau.org  cbatiau@eps.harvard.edu  ISSN 0081-0304

                URL http://www.cfa.harvard.edu/iau/cbat.html

COMET C/2010 G2 (HILL) R. E. Hill reports his discovery of a comet on Catalina Sky Survey CCD

images taken with the 0.68-m Schmidt telescope (discovery observation tabulated below); four co-added 40-s exposures taken in 2′′-3′′ seeing during Apr. 10.45–10.47 UT show a diffuse coma ∼ 6′′–8′′ across with a broad, diffuse tail ∼ 8′′–10′′ long in p.a. 200^{o}. Four co-added 60-s follow-up images by R. A. Kowalski and G. M. Hupe with the Mt. Lemmon 1.5-m reflector on Apr. 10.5 also show the faint, broad, straight tail in p.a. ∼ 200^{o}; additional images with the same telescope by S. M. Larson on Apr. 11.44–11.46 show a 20′′ tail in p.a. 200^{o}. After posting on the Minor Planet Center’s ‘NEOCP’ webpage, other cometary astrometrists have also commented on this object’s cometary appearance during Apr. 11.2–11.7. A. Novichonok and D. Chestnov (remotely with a 36-cm Maksutov-Newtonian telescope at Tzec Maun Observatory near Cloudcroft, NM, U.S.A.) found an obvious faint coma with diameter 0′.2 and no tail. W. H. Ryan and E. V. Ryan (Magdalena Ridge Observatory, 2.4-m reflector) find a coma and a faint tail in p.a. ∼ 200^{o} on R-band images. E. Guido and G. Sostero (remotely using a 0.25-m reflector at the RAS Observatory, Mayhill, NM) report that eighteen co-added 120-s unfiltered exposures reveal a diffuse coma nearly 10′′ in diameter; H. Sato, Ota-ku, Tokyo, Japan (also remotely using a 0.25-m reflector at the same site) measures a 15′′ coma. R. Ligustri, Udine, Italy (remotely using a 0.40-m reflector at the RAS Observatory, Moorook, Australia) writes that eleven 180-s co-added images show a faint 20′′ tail in p.a. 235^{o}.

2010 UT α_{2000} δ_{2000} Mag. Observer
Apr. 10.42936 16^{h}44^{m}05.^{s}35 +10^{o}27′21′′.1 19.2 Hill

The available astrometry, the following very preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2010-G87:

T = 2011 June 19.639 TT  ω = 150.^{o}748
Ω = 244.701 }2000.0
q = 1.25708 AU  i = 101.001

COMETS 234P/LINEAR AND 235P/LINEAR Comet P/2010 E4 = 2002 CF_{140} (cf. IAUC 9126) has been given the

permanent number 234P. Comet P/2010 F2 = P/2002 FA_{9} (cf. IAUC 9130) has been given the permanent number 235P.

2010 April 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2010 CBAT  Daniel W. E. Green